OpenStudy (animallover8):

~Properties in Advanced Euclidean Geometry~ Fan and medal guaranteed for logical, real life examples. Give a geometric example of the Symmetric Property. Give an algebraic example of the Transitive Property. Give an algebraic example of the Reflexive Property. Give a geometric example of the Commutative Property. Give an algebraic example of the Distributive Property. This is for my Advanced Euclidean Geometry - Section 2.10 review. Thank you for your time.

OpenStudy (daniel.ohearn1):

First symmetry implies same shape different orientation, like if we slice a piece of pie in half perfectly..

OpenStudy (daniel.ohearn1):

Transitive is x+y=z and x+y=a therefore what else equals z?

OpenStudy (daniel.ohearn1):

a+b+c+d=e b+c+d+e=12 Transitive Property a+2b+2c+2d=12